# Interacting light thermal-relic dark matter: self-consistent cosmological bounds

### v1

@@ I Introduction
In this work, we present the first self-consistent cosmological likelihood analyses of the effects of DM mass and its interactions using CMB measurements from Planck 2018 [5]. We derive the first joint bounds on the i) DM mass and DM-baryon scattering cross section, ii) DM mass and late-time annihilation cross section, and iii) DM mass, DM-baryon scattering cross section, and late-time annihilation cross section for the specific scenario of DM that possesses magnetic and electric dipole moments. Case iii expands upon previous work, which did not account for the effects of DM mass and did not consider the joint effects of the electric and magnetic dipole interactions [51, 52]. In all cases, we assume that DM achieves its relic abundance by annihilating into SM particles in the early Universe, but we do not enforce that this annihilation at freeze-out is tied to the elastic scattering or late-time annihilation cross section. We find that in the range of thermal-relic masses relevant for BBN ($\lesssim 20$ MeV), the effects of mass and DM interactions can be degenerate and modeling them simultaneously often leads to an improvement in the sensitivity of CMB measurements to DM parameters.
@@ II Thermal-relic DM
We assume that the observed DM abundance is achieved through a thermal freeze-out mechanism [53, 54], and determined by the specific details of the freeze-out process. In this scenario, DM maintains thermal equilibrium through its interactions with SM particles in the early Universe, until the temperature of the thermal bath drops below the mass of DM. However, our analyses we do not require that the freeze-out processes that set the relic abundance are the same as those that are active at later times—deviations from the standard Lee-Weinberg scenario [55] arise, e.g., in models of coannihilation [56], coscattering [57, 58], and semi-annihilation [59].
@@ III.1 DM mass and DM-baryon scattering
In order to compute CMB power spectra in a scenario that simultaneously accounts for the effects of light thermal-relic DM mass and elastic scattering with baryons, we implement the contributions of thermally coupled DM (through its effects on $Y_{\mathrm{p}}$ and $N_{\mathrm{eff}}$) into a modified Boltzmann code CLASS developed for a cosmology that features DM and baryons scattering [42]. In the original version of this code, the value of $Y_{\mathrm{p}}$ is usually treated using standard BBN, where $Y_{\mathrm{p}}$ depends only on baryon density $\Omega_{b}h^{2}$ and extra relativistic species $\Delta N_{\nu}$; and $N_{\mathrm{eff}}$ is a constant parameter, with a default value of 3.044 [60, 61]. In the modified version, we compute $Y_{\mathrm{p}}$ as a function of $\Omega_{b}h^{2}$, $\Delta N_{\nu}$, and $m_{\chi}$, using the publicly-available AlterBBN code [85, 86]; meanwhile, $N_{\mathrm{eff}}$ also depends on the values of $m_{\chi}$. In this work, we do not consider other relativistic degrees of freedom and set $\Delta N_{\nu}=0$.
@@ III.2 DM mass and residual DM annihilation
Residual DM annihilation reduces the CMB temperature power spectrum at in a nearly scale-invariant manner at multipoles $\ell\gtrsim 500$ (green dashed line). Thus, the enhanced power spectrum due to the effects of DM mass (gray dashed line) experiences an overall suppression when incorporating the effects of residual annihilation (blue solid line).
@@ V Summary
We explore DM-baryon with power-law dependence of the scattering cross section on the relative particle velocity, taking into account four values of the power-law index $n\in\{0,2,4,6\}$. We use a modified CLASS code with added capability of simultaneously accounting for effects of light thermal-relic DM mass and the interaction between DM and baryons. The effects of the mass and interaction can add or even cancel out with each other, depending on the values of $n$. We find that, in the range of thermal-relic masses relevant for BBN, the DM mass and cross section are degenerate with each other. Modeling them simultaneously provides more stringent bounds in almost all cases, leading to an improvement in the sensitivity of CMB to the properties of light DM. We then consider a scenario of light thermal-relic DM that features late-time annihilation. We perform the CMB analysis of this joint model and provide improved constraints in the parameter space of DM mass and annihilation cross section. Key results are shown in Figs. 5 and 6.
@@
We expect that the analyses of the CMB measurements with more precision and accuracy at small scales are likely to yield even more stringent constraints on these DM models. Furthermore, including other measurements from structure formation could help break degeneracies between the DM mass and interactions. The same methods we employ in this study may be applied to other thermal-relic DM models probed by CMB anisotropies and other tracers of structure. While for non-thermal relic models, e.g. millicharged DM, entirely new approaches are necessary to seek these candidates. We leave such considerations for future work.

### v2

@@ I Introduction
In this work, we present the first self-consistent cosmological likelihood analyses of the effects of DM mass and its interactions using CMB measurements from Planck 2018 [5]. We derive the first joint bounds on the i) DM mass and DM-baryon scattering cross section, ii) DM mass and late-time annihilation cross section, and iii) DM mass, DM-baryon scattering cross section, and late-time annihilation cross section for the specific scenario of DM that possesses magnetic and electric dipole moments. Case iii expands upon previous work, which did not account for the effects of DM mass and did not consider the joint effects of the electric and magnetic dipole interactions [51, 52, 53]. In all cases, we assume that DM achieves its relic abundance by annihilating into SM particles in the early Universe, but we do not enforce that this annihilation at freeze-out is tied to the elastic scattering or late-time annihilation cross section. We find that in the range of thermal-relic masses relevant for BBN ($\lesssim 20$ MeV), the effects of mass and DM interactions can be degenerate and modeling them simultaneously often leads to an improvement in the sensitivity of CMB measurements to DM parameters.
@@ II Thermal-relic DM
We assume that the observed DM abundance is achieved through a thermal freeze-out mechanism [54, 55], and determined by the specific details of the freeze-out process. In this scenario, DM maintains thermal equilibrium through its interactions with SM particles in the early Universe, until the temperature of the thermal bath drops below the mass of DM. However, in our analyses we do not require that the freeze-out processes that set the relic abundance are the same as those that are active at later times—deviations from the standard Lee-Weinberg scenario [56] arise, e.g., in models of coannihilation [57], coscattering [58, 59], and semi-annihilation [60].
@@ III.1 DM mass and DM-baryon scattering
In order to compute CMB power spectra in a scenario that simultaneously accounts for the effects of light thermal-relic DM mass and elastic scattering with baryons, we implement the contributions of thermally coupled DM (through its effects on $Y_{\mathrm{p}}$ and $N_{\mathrm{eff}}$) into a modified Boltzmann code CLASS developed for a cosmology that features DM and baryon scattering [42]. In the original version of this code, the value of $Y_{\mathrm{p}}$ is usually treated using standard BBN, where $Y_{\mathrm{p}}$ depends only on baryon density $\Omega_{b}h^{2}$ and extra relativistic species $\Delta N_{\nu}$; and $N_{\mathrm{eff}}$ is a constant parameter, with a default value of 3.044 [61, 62, 86]. In order to account for the effects of light thermal-relic DM mass $m_{\chi}$, we first use the publicly-available AlterBBN code [87, 88] to generate new BBN predictions for $Y_{\mathrm{p}}$, as a function of $\Omega_{b}h^{2}$, $\Delta N_{\nu}$, and $m_{\chi}$; and for $N_{\mathrm{eff}}$, as a function of $m_{\chi}$. Then we update the BBN interpolation table and related BBN calculation for $Y_{\mathrm{p}}$ within CLASS code, as well as the calculation of $N_{\mathrm{eff}}$. In this work, we do not consider other relativistic degrees of freedom and set $\Delta N_{\nu}=0$.
@@ III.2 DM mass and residual DM annihilation
Residual DM annihilation reduces the CMB temperature power spectrum in a nearly scale-invariant manner at multipoles $\ell\gtrsim 500$ (green dashed line). Thus, the enhanced power spectrum due to the effects of DM mass (gray dashed line) experiences an overall suppression when incorporating the effects of residual annihilation (blue solid line).
@@ V Summary
We explore DM-baryon scattering with power-law dependence of the scattering cross section on the relative particle velocity, taking into account four values of the power-law index $n\in\{0,2,4,6\}$. We use a modified CLASS code with added capability of simultaneously accounting for effects of light thermal-relic DM mass and the interaction between DM and baryons. The effects of the mass and interaction can add or even cancel out with each other, depending on the values of $n$. We find that, in the range of thermal-relic masses relevant for BBN, the DM mass and cross section are degenerate with each other. Modeling them simultaneously provides more stringent bounds in almost all cases, leading to an improvement in the sensitivity of CMB to the properties of light DM. We then consider a scenario of light thermal-relic DM that features late-time annihilation. We perform the CMB analysis of this joint model and provide improved constraints in the parameter space of DM mass and annihilation cross section. Key results are shown in Figs. 5 and 6.
@@
We expect that the analyses of the CMB measurements with more precision and accuracy at small scales, such as Simons Observatory [95] and CMB-S4 [96, 97], are likely to yield even more stringent constraints on these DM models. Furthermore, including other measurements from structure formation could help break degeneracies between the DM mass and interactions. The same methods we employ in this study may be applied to other thermal-relic DM models probed by CMB anisotropies and other tracers of structure. While for non-thermal relic models, e.g. millicharged DM, entirely new approaches are necessary to seek these candidates. We leave such considerations for future work.